OpenStudy (anonymous):

1. Select the proper equation in point-slope form for the given point and slope. Point: (7, 4); Slope: 6 (Points : 1) y – 7 = 6(x + 4) y – 4 = 6(x – 7) y – 4 = 6(x + 7) y + 4 = 6(x – 7) Question 2. 2. Select the proper equation in point-slope form for the given point and slope. Point: (–2, 4); Slope: 5 (Points : 1) y + 4 = 5x – 10 y + 4 = 5(x + 2) y – 4 = 5(x + 2) y – 4 = 5(x – 2) Question 3. 3. Select the proper equation in point-slope form for the given point and slope. Point: ; Slope: (Points : 1)

OpenStudy (anonymous):

please help will return the favor

OpenStudy (anonymous):

i need to know if my answers is correct is 1b 2c

OpenStudy (anonymous):

point slope form \[(y-y_1) = m(x-x_1)\] (x1,y1) = given points m = slope (given)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

Your answers are correct.

OpenStudy (anonymous):

both of them????

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

can you help me with just FEW more

OpenStudy (anonymous):

please??

OpenStudy (anonymous):

sure.

OpenStudy (anonymous):

thanks so much ....elect the proper equation in point-slope form for the given point and slope. Point: ; Slope:

OpenStudy (anonymous):

the ponits are 2.-6 slops is -3/4

OpenStudy (anonymous):

what will be your answer?

OpenStudy (anonymous):

y-6=-3/4x+2 y-6=3\4x-2 y+6=-3/4x-2 y+6=-3/4x+2

OpenStudy (anonymous):

@vickielt you know the answer

OpenStudy (anonymous):

i thinks it the 3rd one

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

A yearbook printer charges based on the number of pages printed. Here is a table that shows the cost of some recent yearbooks. Number of pages 50 100 150 200 Cost 1.55 2.25 2.95 3.65 Which equation gives the cost, y, in terms of the number of pages, x? (Points : 1) y = 0.014x + 49.15 y = 0.014x + 1.55 y = 0.014x + 0.85 y = 58.8x + 0.85

OpenStudy (anonymous):

Find slope m = (y2-y1)/(x2-x1) (x1,y1), (x2,y2) are given in table ( x = number of pages, y = cost) then use point slope formula (y-y1)= m(x-x1)

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

which part?

OpenStudy (anonymous):

everything but i think the answers is b

OpenStudy (anonymous):

m = (change in price)/(change in no of pages) m = (2.25 - 1.55)/(100-50) m = 0.70/50 = 0.014 m = change of rate per page x = number of pages y = cost using point slope formula we can find our equation (y-y1)= m(x-x1) and your answer is correct

OpenStudy (anonymous):

are you sure becuase i have been gussing are you just telling things os i can fell

OpenStudy (anonymous):

I am 100 % sure. If you want you can ask someone else for your surety.

OpenStudy (anonymous):

im not trying to be mean or anything

OpenStudy (anonymous):

Its okay. I understand. :) I would still say if you want to be 100% percent sure then you can ask someone for help

OpenStudy (anonymous):

no one will talk t me

OpenStudy (anonymous):

Wait I will ask someone

OpenStudy (anonymous):

@TheSmartOne can you help?